OpenStudy (anonymous):

Find the area of the triangle 1 1/4 yards and height 4/7 yard but I have the answer I just need to see if its correct? Help

OpenStudy (anonymous):

Is it the base 1 1/4 yards?

OpenStudy (anonymous):

yes

OpenStudy (amorfide):

0.5(bxh) b being base h being height so you do 1 1/4 x 4/7 this gives 20/28 so you get 5/7 then multiply by 1/2 5/14

OpenStudy (anonymous):



OpenStudy (anonymous):

ok this how i got mine 1 1/4x 4/7= 5/4x4/7=20/11

OpenStudy (anonymous):

than i did 20/11x1/2 So where did i went wrong

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

It looks so ugly, but you got my point, right?

OpenStudy (anonymous):

yeah thanks lol